These were accepted as a reasonable account of the last meeting.
RAT will need to give a bit of thought as to exactly what's required here, but something certainly will be. It might be worth talking to Angus.
Still pending ECA's new server room. Defer.
Still quite a few to be done, owned by RAT, US and Services. "Who" to be amended accordingly.
Actions added or revived:
Blog articles suggested:
Report from Computing Executive Group
It's currently looking likely that there'll be three batches of moves:
All dates still subject to change, of course!
Reports from Units
The kerberos master will be moved to one of the slaves tomorrow
morning, first thing. A new disc should then be obtained from Dell for
bevan while it's still in warranty, for the benefit of its
next user. It's a Dell R320.
localhome and apacheconf might be affected by the autofs change
Please try it, and report any oddities.
Dell asked for a SupportAssist log.
It's currently only in the dice layer. Stephen reckoned it would make sense for it to be in the lcfg layer. He'll arrange for that to happen.
One of the drives in the tape library failed. The Spectra support people suggested unplugging it and replugging it, which Neil did. It's working again.
US's wiki page is here.
Topics for discussion
The next meeting will be on Wednesday 14th June at 10:00, in IF-4.31, chaired by Alison.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh